Jane works due north of her home. Her husband Juan works due east. They leave for work at the same time. By the time Jane is 7 miles from home, the distance between then is only 1 mile more than Juans distance from home. How far from home is Juan?
Answer by lwsshak3(11628)

Jane works due north of her home. Her husband Juan works due east. They leave for work at the same time. By the time Jane is 7 miles from home, the distance between them is only 1 mile more than Juans distance from home. How far from home is Juan?
***
let x=Juan's distance from home
7=Jane's distance from home
x+1=distance between them
x^2=(x+1)^2-7^2)=x^2+2x+1-49
2x-48=0
x=24
Juan's distance from home=24 miles